# FreeTacMan: Robot-free Visuo-Tactile Data Collection System for Contact-rich Manipulation

### v1

@@ Acknowledgments
We gratefully acknowledge Huijie Wang for developing the demonstration page, and Zherui Qiu for assisting with the supervision of the user study. We also thank Yixuan Pan, Qingwen Bu, Zhuoheng Li, Jisong Cai, Yuxiang Lu, and Ningbin Zhang for their valuable insights and constructive discussions. Our appreciation also goes to Zhirui Zhang, Xinyu Yang, Fengjie Shen, Taoyuan Huang, and Lekai Chen for their assistance during the experiments. Finally, we extend our sincere gratitude to all user study participants for their thoughtful feedback.
Supplementary Material
@@ Appendix A Hardware Design
| Supplement to Sec. 3.1 in the Main paper.
@@
Seamless Integration into Robotic Platforms. Fig. 5 illustrates the universal gripper interface with quick-swap mounts compatible with both the Piper and Franka arms, as well as the camera scaffold designed for precise alignment with the wrist-mounted camera to ensure consistent perspective. These components demonstrate the plug-and-play modularity of FreeTacMan, enabling seamless integration across diverse robotic platforms without requiring hardware-specific modifications.
@@ Appendix B Data Processing
| Supplement to Sec. 3.2 in the Main paper.
@@ Appendix C Training Details
| Supplement to Sec. 3.3 in the Main paper.
@@
Table 4: Hyperparameters for tactile pretraining.
@@
Details on Tactile Pretraining. We pretrain the tactile encoder $f_{t}$ and projection head $g_{v}$, $g_{t}$ using a CLIP-style contrastive loss with multi-positive sampling. Optimization is performed using AdamW with cosine-annealed learning rates, updating only the tactile encoder $f_{t}$, visual $g_{v}$, and tactile projection $g_{t}$ while keeping visual encoder $f_{v}$ fixed. We train on the sequence of time-aligned visuo-tactile frames, drawing primary positives from the same timestep and secondary positives from the next frame (with wrap-around). Negatives are sampled from a memory bank of size $4096$. Table 4 shows the hyperparameters of pretraining.
Details on Policy Learning. We adopt the action chunking transformer (ACT) [11] architecture to learn visuo-tactile manipulation policies. Both visual and tactile images are  $\,640\times 480$ pixels and fed as stacked inputs to the ACT backbones. At each timestep, the model outputs an action chunk of length $\mathcal{T}$, where each action in the chunk specifies a 6-DOF joint action for robot arm plus a 1-DOF gripper action. To integrate our visuo-tactile pipeline, we add our pretrained tactile encoder $f_{t}$ to extract tactile features. The tactile features are concatenated with visual features and input into the transformer encoder of ACT. Table 5 shows the hyperparameters of policy learning.
Table 5: Hyperparameters for policy learning.
@@
Learning rate $54e-$ 5
Weight decay $41e-$ 4
@@ Appendix D Experimental Setup
| Supplement to Sec. 4.1 in the Main paper.
@@
  • Fragile Cup Manipulation. In this task, the robot is required to grasp a small plastic cup from a rack and place it into a tray on the rack. A trial is considered successful if the cup is stably placed within the tray without any damage or visible deformation to the cup. This task is designed to evaluate the role of tactile feedback in delicate object handling.
  • USB Plugging. The robot must insert a pre-grasped USB into a specified socket on a power strip for this task. A success is recorded if the USB is inserted to a sufficient depth such that it remains securely connected. Tactile information serves as an additional signal to assist with fine alignment.
@@
  • Stamp Pressing. In the stamping task, the robot needs to press a pre-grasped circular stamp onto a sheet of paper laid flat on the table. The task is deemed successful if a clear and complete imprint of the stamp is produced on the paper surface. The stamping task leverages tactile feedback to control contact force during imprinting.
  • Calligraphy Writing. For the calligraphy brush writing task, the robot must use a pre-grasped calligraphy brush to trace the digit "5" following a guide labeled "202" on a piece of cloth. A success is determined if the resulting written pattern can be visually recognized and interpreted as the number "5". In this task, tactile information helps the robot estimate the in-hand brush pose during fine motion.
  • Toothpaste Extrusion. This task requires the robot to extrude toothpaste from a pre-grasped tube onto a sheet of paper placed on the table. The task is considered successful if the toothpaste is deposited on the paper in an amount consistent with typical daily use for brushing. Tactile feedback helps the robot modulate the applied force, ensuring the right amount of toothpaste is dispensed.
  • Tissue Grasping. The robot is tasked with extracting a single tissue from a box. A trial is considered successful if the tissue is pulled out without slipping. Tactile sensing provides real-time feedback to confirm initial contact and ensure a stable grasp throughout the pulling motion.
  • Potato Chip Grasping. In this task, a fragile potato chip—slightly larger in diameter than a plastic cup—is placed atop the cup. The robot must grasp the chip and transfer it to a nearby dish without touching the cup or breaking the chip. A trial is deemed successful if the chip is placed intact into the dish. Tactile sensing plays a critical role in confirming secure contact with the chip while avoiding excessive force that could lead to breakage.
Implementation Details. We conduct all experiments on a PIPER 6-DOF light-weight robotic arm (AGILE·X Robotics) equipped with our FreeTacMan mounted at the wrist, as shown in Fig. 7. The robot is connected to a workstation running Ubuntu 20.04 and ROS Noetic on an NVIDIA RTX 4090 GPU. RGB frames are captured at 30 Hz by HBVCAM fisheye camera and published via ROS topics, and tactile frames from our camera-based tactile sensor are sampled at 30 Hz. Our imitation policy performs inference on the GPU, with policy latency averaging under 20 ms per cycle.
@@
  • Comfort and Ease of Collection: After completing all eight tasks with each of the three systems, participants rated how physically comfortable it was to perform the demonstrations on a scale from 1 (very uncomfortable) to 5 (very comfortable). In the same survey, they scored the ease of collection from 1 (very difficult) to 5 (very easy).
@@
All metric values are normalized, with higher scores indicating better performance, as visualized by the radar chart shown in Fig. 4(e) in the main paper.
@@ Appendix E Additional Experiments
Evaluation on User Study. Table 6 reports completion rate, task duration, slip count, damage count and CPUT for each method across our eight contact-rich tasks.
Table 6: Detailed results of user study. FreeTacMan outperforms existing data collection methods across seven manipulation tasks. Completion Rate is the fraction of successful demonstrations; Time is the average collection duration; Slip Count and Damage Count record object slips and damages during collection; Performance is an overall metric combining efficiency and reliability.
@@
Fragile cup ALOHA 0.4997 11.48 2 14 0.0435
UMI 1.0000 4.26 0 0 0.2348
Ours 1.0000 3.49 0 0 0.2863
USB Plug ALOHA 0.5997 12.40 9 0 0.0484
UMI 0.1665 5.63 27 0 0.0296
Ours 0.9666 4.26 2 0 0.2270
Calligraphy Writing ALOHA 0.6664 14.44 2 0 0.0461
UMI 1.0000 4.72 2 0 0.2120
Ours 1.0000 3.93 0 0 0.2543
Stamp Pressing ALOHA 0.7330 8.18 3 2 0.0896
UMI 0.4997 4.05 26 0 0.1235
Ours 1.0000 2.86 1 0 0.3497
Toothpaste Dispensing ALOHA 0.5665 8.73 16 0 0.0649
UMI 0.1998 6.41 6 0 0.0312
Ours 0.9332 5.34 1 0 0.1746
Tissue Grasping ALOHA 0.8665 6.44 5 0 0.1346
UMI 0.9332 3.29 3 0 0.2837
Ours 1.0000 2.44 0 0 0.4103
Chip Grasping ALOHA 0.5330 12.94 5 4 0.0412
UMI 1.0000 4.06 0 0 0.2464
Ours 0.9666 3.69 1 0 0.2617
  • Completion Rate: Across all tasks, FreeTacMan achieves the highest completion rates, reaching 100% (Texture class, Fragile cup, Stamp pressing, Tissue grasping, Calligraphy) or near-perfect rates (USB plugging, Chip grasping, Toothpaste dispensing). In contrast, ALOHA’s rates range from 49.97% (Fragile cup) to 86.65% (Tissue grasping), and UMI’s range from 16.65% (USB plug) to 100% (Calligraphy, Chip grasping). This demonstrates that real-time tactile feedback enhances the human operator’s ability to execute and complete precise, contact-rich tasks.
  • Task Duration: FreeTacMan demonstrates the shortest task durations, averaging between 2.44s (Tissue grasping) and 5.34s (Toothpaste dispending). UMI’s times range from 3.29s to 6.41s, while ALOHA’s are substantially higher (6.44s to 14.44s). Handheld paradigm outperforms ALOHA in data collection efficiency due to its intuitive feedback, while in-situ feedback of FreeTacMan leads to faster data collection.
  • Slip and Damage Count: High slip counts indicate poor force and slippage feedback. ALOHA records up to 27 slips (USB insertion) and 14 damages (Fragile cup), reflecting its indirect force perception. UMI improves the force feedback and records zero damage events, yet its multi-link transmission still permits noticeable slippage in force-sensitive tasks (e.g., 26 slips during Stamp pressing). In contrast, FreeTacMan limits slips to at most 2 and records zero damage across all tasks, evidencing its high-fidelity, low-latency tactile feedback.
  • CPUT: As a combined metric of efficiency and reliability, CPUT highlights the overall advantage of FreeTacMan: it achieves scores from 0.1746 (Toothpaste dispensing) to 0.4103 (Tissue grasping), substantially outperforming ALOHA and UMI. For tasks like Stamp pressing, the CPUT of FreeTacMan reaches 0.3497, which exceeds ALOHA and UMI by nearly 3 times, confirming its advantage under precise contact requirements.
@@
Figure 8: Visual Attention Heatmaps in Texture Classification. Cross-attention overlays on the visual input: Left two panels (Vision-only ACT): (a) before grasp and (b) after grasp. Right two panels (Visuo-tactile ACT): (c) before grasp and (d) after grasp.
@@
Fig. 8 shows the decoder cross-attention overlays on the visual input during the texture classification task. In the first row, we show two snapshots from the vision-only ACT model: the left image is at the initial state of inference, where attention is correctly focused on the target cylindrical object; the right image is after the gripper has lifted the object, the model’s attention is diffusely allocated to both the red and blue bins due to the lack of tactile cues. In the second row, we show the corresponding stages for our model with tactile integration: initial attention in the left image is on the target object again, but after lift (right), the model uses tactile feedback to recognize texture and concentrates its attention exclusively on the correct (blue) bin. This comparison demonstrates how tactile information refines the decision of policy during inference.
Fig. 9 depicts how the ACT model’s decoder cross-attention shifts over the tactile image during inference of texture classification. The left image shows the attention map just before the gripper makes contact: since no deformation has happened, attention is broadly distributed over the gel layer. In contrast, the right image captures the moment immediately after contact, where the model concentrates its attention on the region of deformation that corresponds to the object’s surface texture. This dynamic reallocation of attention enables the policy to detect fine-grained tactile cues and infer both texture and object geometry in real time.

### v2

@@ VI Acknowledgments
We gratefully acknowledge Huijie Wang for helping to develop the demonstration page, and Zherui Qiu for assisting with the supervision of the user study. We also thank Yixuan Pan, Qingwen Bu, Zhuoheng Li, Jisong Cai, Yuxiang Lu, and Ningbin Zhang for their valuable insights and constructive discussions. Our appreciation also goes to Yingxin Wang, Zhirui Zhang, Xinyu Yang, Fengjie Shen, Taoyuan Huang, and Lekai Chen for their assistance during the data collection. Finally, we extend our sincere gratitude to JAKA for their generous support in the collection of our dataset.
@@ Appendix A Hardware Design
| Supplement to Sec.˜III-A in the Main paper.
@@
Seamless integration into robotic platforms. Fig.˜9 illustrates the universal gripper interface with quick-swap mounts compatible with both the Piper and Franka arms, as well as the camera scaffold designed for precise alignment with the wrist-mounted camera to ensure consistent perspective. These components demonstrate the plug-and-play modularity of FreeTacMan, enabling seamless integration across diverse robotic platforms without requiring hardware-specific modifications.
@@ Appendix B Data Processing
| Supplement to Sec.˜III-B in the Main paper.
@@ Appendix C Training Details
| Supplement to Sec.˜III-C in the Main paper.
@@
TABLE V: Hyperparameters for tactile pretraining.
@@
Details on tactile pretraining. We pretrain the tactile encoder $f_{t}$ and projection head $g_{v}$, $g_{t}$ using a CLIP-style contrastive loss with multi-positive sampling. Optimization is performed using AdamW with cosine-annealed learning rates, updating only the tactile encoder $f_{t}$, visual $g_{v}$, and tactile projection $g_{t}$ while keeping visual encoder $f_{v}$ fixed. We train on the sequence of time-aligned visuo-tactile frames, drawing primary positives from the same timestep and secondary positives from the next frame (with wrap-around). Negatives are sampled from a memory bank of size $4096$. Table V shows the hyperparameters of pretraining.
Details on policy learning. We adopt the action chunking transformer (ACT) [26] architecture to learn visuo-tactile manipulation policies. Both visual and tactile images are  $640\times 480$ pixels and fed as stacked inputs to the ACT backbones. At each timestep, the model outputs an action chunk of length $\mathcal{T}$, where each action in the chunk specifies a 6-DOF joint action for robot arm plus a 1-DOF gripper action. To integrate our visuo-tactile pipeline, we add our pretrained tactile encoder $f_{t}$ to extract tactile features. The tactile features are concatenated with visual features and input into the transformer encoder of ACT. Table˜VI shows the hyperparameters of policy learning.
TABLE VI: Hyperparameters for policy learning.
@@
Learning rate 4e54e-5
Weight decay 1e41e-4
@@ Appendix D Experimental Setup
| Supplement to Sec.˜IV-B in the Main paper.
@@
  • Fragile cup. In this task, the robot is required to grasp a fragile cup from a rack and place it into a tray on the rack. A trial is considered successful if the cup is stably placed within the tray without any damage or visible deformation to the cup. This task is designed to evaluate the role of tactile feedback in delicate object handling.
  • USB plug. The robot must insert a pre-grasped USB into a specified socket on a power strip for this task. A success is recorded if the USB is inserted to a sufficient depth such that it remains securely connected. Tactile information serves as an additional signal to assist with fine alignment.
@@
  • Stamp press. In the stamping task, the robot needs to use a pre-grasped, spring-loaded stamp to produce a clear and complete imprint onto a sheet of paper laid flat on the table. The task is deemed successful if a clear and complete imprint of the stamp is produced on the paper surface. The stamping task leverages tactile feedback to control contact force during imprinting.
  • Calligraphy. For the calligraphy brush writing task, the robot must use a pre-grasped calligraphy brush to trace the digit "5" following a guide labeled "202" on a piece of cloth. A success is determined if the resulting written pattern can be visually recognized and interpreted as the number "5". During this task, tactile information assists the robot in both estimating the in-hand brush pose for fine motion control and ensuring the brush-lifting action between the first and the second stroke.
Implementation details. We conduct all experiments on a PIPER 6-DOF light-weight robotic arm (AGILE·X Robotics) equipped with our FreeTacMan mounted at the wrist, as shown in Fig.˜11. The robot is connected to a workstation running Ubuntu 20.04 and ROS Noetic on an NVIDIA RTX 4090 GPU. RGB frames are captured at 30 Hz by HBVCAM fisheye camera and published via ROS topics, and tactile frames from our camera-based tactile sensor are sampled at 30 Hz. Our imitation policy performs inference on the GPU, with policy latency averaging under 20 ms per cycle.
@@
  • Comfort and ease of collection: After completing five tasks with each of the three systems, participants rated how physically comfortable it was to perform the demonstrations on a scale from 1 (very uncomfortable) to 5 (very comfortable). In the same survey, they scored the ease of collection from 1 (very difficult) to 5 (very easy).
@@
All metric values are normalized, with higher scores indicating better performance, as visualized by the radar chart shown in Fig. 6(d) in the main paper.
@@ Appendix E Additional Experiments
Evaluation on user study. Table VII reports completion rate, task duration, slip count, damage count and CPUT for each method across five contact-rich tasks.
TABLE VII: Detailed results of user study. FreeTacMan outperforms existing data collection methods across five manipulation tasks. Completion Rate is the fraction of successful demonstrations; Time is the average collection duration; Slip Count and Damage Count record object slips and damages during collection; Performance is an overall metric combining efficiency and reliability.
@@
Fragile Cup ALOHA 0.5274 11.19 2 14 0.0471
UMI 1.0000 4.19 0 0 0.2386
Ours 1.0000 3.50 0 0 0.2854
USB Plug ALOHA 0.6108 12.63 9 0 0.0483
UMI 0.2220 5.55 27 0 0.0400
Ours 0.9722 4.24 2 0 0.2292
Calligraphy ALOHA 0.5553 14.89 2 0 0.0373
UMI 1.0000 5.03 2 0 0.1989
Ours 1.0000 4.17 0 0 0.2401
Stamp Press ALOHA 0.7775 8.09 3 2 0.0962
UMI 0.5553 3.91 26 0 0.1421
Ours 1.0000 2.98 1 0 0.3356
  • Completion rate: Across all tasks, FreeTacMan achieves the highest completion rates, reaching 100% (Texture classification, Fragile cup, Stamp press, Calligraphy) or near-perfect rates (USB plug). In contrast, ALOHA’s rates range from 52.74% (Fragile cup) to 86.11% (Texture classification), and UMI’s range from 22.20% (USB plug) to 100% (Texture classification, Fragile cup, Calligraphy). This demonstrates that real-time tactile feedback enhances the human operator’s ability to execute and complete precise, contact-rich tasks.
  • Task duration: FreeTacMan demonstrates the shortest task durations, averaging between 2.98s (Stamp press) and 4.24s (USB plug). UMI’s times range from 3.27s to 5.55s, while ALOHA’s are substantially higher (8.09s to 14.89s). Handheld paradigm outperforms ALOHA in data collection efficiency due to its intuitive feedback, while in-situ feedback of FreeTacMan leads to faster data collection.
  • Slip and damage count: High slip counts indicate poor force and slippage feedback. ALOHA records up to 27 slips (USB plug) and 14 damages (Fragile cup), reflecting its indirect force perception. UMI improves the force feedback and records zero damage events, yet its multi-link transmission still permits noticeable slippage in force-sensitive tasks (e.g., 26 slips during Stamp press). In contrast, FreeTacMan limits slips to at most 2 and records zero damage across all tasks, evidencing its high-fidelity, low-latency tactile feedback.
  • CPUT: As a combined metric of efficiency and reliability, CPUT highlights the overall advantage of FreeTacMan: it achieves scores from 0.2292 (USB plug) to 0.3356 (Stamp press), substantially outperforming ALOHA and UMI. For tasks like stamp press, the CPUT of FreeTacMan reaches 0.3497, which exceeds ALOHA and UMI by nearly 3 times, confirming its advantage under precise contact requirements.
@@
Figure 12: Visual attention heatmaps in texture classification. Cross‐attention overlays on the visual input: Left two panels (Vision‐only ACT): (a) before grasp and (b) after grasp. Right two panels (Visuo‐tactile ACT): (c) before grasp and (d) after grasp.
@@
Fig.˜12 shows the decoder cross-attention overlays on the visual input during the texture classification task. In the first row, we show two snapshots from the vision-only ACT model: the left image is at the initial state of inference, where attention is correctly focused on the target cylindrical object; the right image is after the gripper has lifted the object, the model’s attention is diffusely allocated to both the red and blue bins due to the lack of tactile cues. In the second row, we show the corresponding stages for our model with tactile integration: initial attention in the left image is on the target object again, but after lift (right), the model uses tactile feedback to recognize texture and concentrates its attention exclusively on the correct (blue) bin. This comparison demonstrates how tactile information refines the decision of policy during inference.
Fig.˜13 depicts how the ACT model’s decoder cross-attention shifts over the tactile image during inference of texture classification. The left image shows the attention map just before the gripper makes contact: since no deformation has happened, attention is broadly distributed over the gel layer. In contrast, the right image captures the moment immediately after contact, where the model concentrates its attention on the region of deformation that corresponds to the object’s surface texture. This dynamic reallocation of attention enables the policy to detect fine-grained tactile cues and infer both texture and object geometry in real time.